A rectangle has twice the area of a square. The length of the rectangle is 12 cm greater and the width is 8 cm greater than a side of a square . Find the perimeter of the square
Answer by mananth(16946)

side of square = x
Are of square = x^2
Rectangle length = (x+12)
rectangle width = (x+8)
Area of rectangle = (x+12)(x+8)
(x+12)(x+8) =2x^2
x^2+20x+96=2x^2
x^2-20x-96=0
x^2-24x+4x-96=0
x(x-24)+4(x-24)=0
(x-24)(x+4)=0
ignore negative
x= 24 the side of square
Perimeter of square = 24 *4 = 96cm
m.ananth@hotmail.ca